# Instance Generation for Meta-Black-Box Optimization through Latent Space Reverse Engineering

### v1

@@ Abstract
To relieve intensive human-expertise required to design optimization algorithms, recent Meta-Black-Box Optimization (MetaBBO) researches leverage generalization strength of meta-learning to train neural network-based algorithm design policies over a predefined training problem set, which automates the adaptability of the low-level optimizers on unseen problem instances. Currently, a common training problem set choice in existing MetaBBOs is well-known benchmark suites CoCo-BBOB. Although such choice facilitates the MetaBBO’s development, problem instances in CoCo-BBOB are more or less limited in diversity, raising the risk of overfitting of MetaBBOs, which might further results in poor generalization. In this paper, we propose an instance generation approach, termed as LSRE, which could generate diverse training problem instances for MetaBBOs to learn more generalizable policies. LSRE first trains an autoencoder which maps high-dimensional problem features into a 2-dimensional latent space. Uniform-grid sampling in this latent space leads to hidden representations of problem instances with sufficient diversity. By leveraging a genetic-programming approach to search function formulas with minimal L2-distance to these hidden representations, LSRE reverse engineers a diversified problem set, termed as Diverse-BBO. We validate the effectiveness of LSRE by training various MetaBBOs on Diverse-BBO and observe their generalization performances on either synthetic or realistic scenarios. Extensive experimental results underscore the superiority of Diverse-BBO to existing training set choices in MetaBBOs. Further ablation studies not only demonstrate the effectiveness of design choices in LSRE, but also reveal interesting insights on instance diversity and MetaBBO’s generalization.
@@ 1 Introduction
Black-Box Optimization (BBO) problems, with neither the objective function nor corresponding derivative is accessible, are challenging to solve. Although Evolutionary Computation (EC) methods, such as Genetic Algorithm (GA) (Holland 1992), Particle Swarm Optimization (PSO) (Kennedy and Eberhart 1995), Differential Evolution (DE) (Storn and Price 1997), as well as their modern (self-)adaptive variants are widely recognized as effective solutions, they often rely on labor-intensive tuning or even re-design for novel scenario adaption.
@@
To automate such manual algorithm design tasks, recent Meta-Black-Box Optimization (MetaBBO) researches propose learning generalizable design policy through meta-learning (Ma et al. 2023, 2025d). MetaBBO naturally forms a bi-level algorithmic structure. At the meta level, a learnable algorithm design policy (e.g., a neural network) outputs flexible design choices (e.g., parameter control (Ma et al. 2024a) and algorithm selection (Guo et al. 2024)) for low-level optimizer. At the lower level, a BBO optimizer is deployed to optimize target problem with the given design choice. The meta-objective is to train a design policy that enhances low-level BBO process.
@@
Given that MetaBBO essentially follows meta-learning paradigm (Thrun and Pratt 1998; Finn, Abbeel, and Levine 2017), it unavoidably requires domain generalization designs (Li et al. 2018), i.e., meta-augmentation of training data (Rajendran, Irpan, and Jang 2020), so as to ensure the adaption during the testing time. To be specific, as shown in the yellow box of Fig. 1, existing MetaBBO approaches typically train their policies over a training problem set. The diversity of this set assures that the meta-level policy generalizes across diverse optimization problem characteristics (Ma et al. 2025b). Common choices of the problem set in existing MetaBBO approaches are synthetic benchmarks such as CoCo-BBOB (Hansen et al. 2010a). Although these representative problem instances hold a long-standing history and reputation in the evaluation of BBO methods, they are primarily hand-crafted to differentiate algorithms apart from each other in terms of optimization performance (Langdon and Poli 2007). Recent studies (Ma et al. 2025d, c) reveal that such design bias may not keeps pace with the training&evaluation objectives of MetaBBO, i.e., generalization toward unseen problems. One possible solution is deriving realistic instances directly from real-world application (Kumar et al. 2020). However, collecting and formulating realistic problems require deep expertise. More importantly, the corresponding expensive simulation costs makes this solution impractical (Muñoz and Smith-Miles 2020). A more promising way is generating synthetic problem under proper diversity measurement (Tian et al. 2020; Vermetten et al. 2025), however, to the best of our knowledge, related researches are still very limited.
@@
To address the aforementioned limitations, we propose a novel generation framework that discovers diverse synthetic optimization problems through latent space reverse engineering, termed as LSRE, to advance the benchmark development in BBO and MetaBBO fields. Specifically, LSRE comprises following key designs: 1) an instance space analysis (Smith-Miles and Muñoz 2023) method that trains an autoencoder to map the Exploratory Landscape Analysis (ELA) (Mersmann et al. 2011) features of problem instances into 2-dimensional latent space; 2) an expressive symbol set including operands and operators that spans a comprehensive functional instance space; 3) a genetic programming method that effectively searches for function instance within the latent space, with an enhanced local search strategy and parallel acceleration. By grid-sampling 256 instance points from the latent space, we repeatedly call LSRE to search optimal function formula for each of them, resulting in a collection of 256 synthetic problem instances Diverse-BBO with adequate diversity. We validate the diversity advantage of our method against existing baseline benchmarks through large-scale MetaBBO benchmarking. The comprehensive experimental results demonstrate the effectiveness of our LSRE framework and the superiority of the generated Diverse-BBO to baseline benchmarks. We now summarize the core contributions of this paper:
@@ 2 Related Works
Existing MetaBBO approaches can be divided into several lines according to their specific algorithm design tasks: 1) Algorithm Selection (AS), where predefined algorithms/operators are selected by the meta-level policy to solve given problems. Early researches explored selecting operators within DE frameworks (Sharma et al. 2019; Tan and Li 2021), and more recent approaches have investigated dynamic switching between multiple advanced optimizers along the optimization process (Guo et al. 2024); 2) Algorithm Configuration (AC), where given an existing optimizer, the meta-level policy dictates desired parameter values dynamically along the optimization process (Sun et al. 2021; Lange et al. 2023; Ma et al. 2024a; Guo et al. 2025b); 3) Algorithm Generation (AG), where the meta-level policy constructs both individual components and the entire workflow to synthesize novel BBO algorithms through algorithm workflow composition (Zhao et al. 2024), novel update rules generation (Chen et al. 2024) and language modelling (Ma et al. 2024b; Liu et al. 2024; van Stein and Bäck 2024). In addition, sutdies on MetaBBO have also been extended to automated feature learning (Ma et al. 2025b; Guo et al. 2025a), efficient offline learning (Ma et al. 2025a, e) and direct solution manipulation (Li et al. 2024, 2025).
@@
Existing MetaBBO methods typically adopt synthetic problems from well-established BBO benchmarks (eg., CoCo-BBOB (Hansen et al. 2010a), IOHProfiler (Doerr et al. 2019), IEEE CEC Competition series (Mohamed et al. 2021)) and platforms (eg., MetaBox-v1/v2 (Ma et al. 2023, 2025c)). While these benchmark functions have been widely recognized and used for evaluating BBO methods, they are primarily hand-crafted and are collected with inherent biases (Hooker 1995), aimed at creating performance distinctions between optimization algorithms (Langdon and Poli 2007). The existing benchmarks’ design bias and lack of diversity may constrain the MetaBBO to learn generalizable design policies that excel at not only the training problem distribution but also unseen problems with various landscape features (Ma et al. 2025d, b). Constructing and collecting realistic problems (Kumar et al. 2020; Hwang et al. 2010) could be a potential solution, but it requires deep expertise and is further limited by simulation costs (Muñoz and Smith-Miles 2020), ultimately rendering the solution impractical.
Generating diverse synthetic problems might be more promising, yet several challenges persist: 1) The random function generator (Tian et al. 2020) lacks control over the characteristics and diversity of generated problems; 2) A recent line of works explore the possibility of augmenting new problem instances through affine combination of existing instances in BBO benchmark (Dietrich and Mersmann 2022). A representative method is MA-BBOB (Vermetten et al. 2025), which claimed achieved efficient and fine-grained problem generation. However, a key limitation is such approaches might restrict themselves within the selected composition functions. 3) In the light of symbolic regression (Schmidt and Lipson 2009), some exploratory studies leveraged symbolic discovery tool such as Genetic Programming (GP) (Koza 1992) to generate diversified instances by searching corresponding mathematical symbolic formulas (Muñoz and Smith-Miles 2020; Long et al. 2023). However, such researches are still limited. More importantly, given the inherent expensive running of GP and the design bottlenecks in these methods, their actual generation effectiveness is still unsatisfactory.
In this paper, we propose LSRE to address the limitations in the third line of work in both effectiveness and efficiency side. On the one hand, we incorporate LSRE with several novel designs: 1) Effective latent instance space analysis through autoencoder; 2) Expressive symbol set enabling both efficient generation and flexible operations; 3) Cross-dimensional local search strategy to further enhance the GP searching. On the other hand, we introduce high-performance distributed acceleration techniques to facilitate efficient GP running.
@@ 3.1 Latent Instance Space Analysis
To align with practical usage of potential users, one of the most important issues in our framework is to pre-define a target problem range $\mathcal{D}$ as the basis for constructing corresponding the latent instance space. While interests of different users may vary, in this paper we primarily select CoCo-BBOB (Hansen et al. 2010a) to serve as $\mathcal{D}$ due to its ease of acquisition and long-standing reputation within BBO evaluation domain. We have to note that our framework can be easily extended to other user-specific problem ranges. We leave step-by-step instructions for such cases at Appendix A.1. Given the 24 basic functions ($f_{1}\sim f_{24}$) in CoCo-BBOB, we first augment these functions with five different problem dimensions $[2,5,10,30,50]$, resulting in 120 functions. Next, we further generate a sufficient number of instances by applying random rotation and shift transformations to these 120 functions:
@@
We construct latent instance space of $\mathcal{D}$ by combining the Exploratory Landscape Analysis (ELA) (Mersmann et al. 2011) and popular dimension reduction strategy autoencoder (Rumelhart, Hinton, and Williams 1985). We illustrate the workflow of such latent instance space analysis in Fig. 2, and briefly introduce how these two techniques cooperate with each other:
ELA. It is a widely used tool for instance space analysis (Škvorc, Eftimov, and Korošec 2020) of BBO problems. Groups of analytic or statistical features, such as convexity, global structure and meta-model fitting are computed in ELA, which has been proven effective in the algorithm selection domain (Kerschke and Trautmann 2019). Given a BBO problem instance $f\in\mathcal{D}$, we obtain its ELA features vector $E_{f}$ using an open-source ELA computation package (Prager and Trautmann 2023): pflacco [^1] https://github.com/Reiyan/pflacco. Specifically, we select a subset of representative and computationally efficient features as $E_{f}$. The selection standard and concrete feature list are provided in Appendix A.2 .
@@
We leave the settings of the autoencoder such as its architecture and corresponding training protocol in Appendix A.3. Once the training of the autoencoder finishes, we could obtain latent representation $h\in\mathbb{R}^{2}$ of a BBO problem instance through the encoder. More importantly, the well-trained model allows us to measure the diversity of problem instances in a 2-dimensional space, which further facilitates uniform instance sampling from this hidden space, since directly sampling from high-dimensional ELA space is neither efficient nor practical. It is worth mentioning that principal component analysis (PCA) (Jolliffe 2011) is also a popular dimension reduction approach widely adopted in machine learning community. The reason behind our autoencoder choice is that: 1) PCA shows effectiveness in cases where features hold linear correlation assumption (Almotiri, Elleithy, and Elleithy 2017; Wang, Yao, and Zhao 2016), however, different feature groups in ELA are independently computed; 2) We have conducted a preliminary study on using PCA for ELA feature reduction, the results show that the first two principle components of certain portion of instances in $\mathcal{D}$ can not surpass 51% importance threshold. We also provide an ablation study in Sec. 4.3.
@@ 3.2 Genetic Programming Search
Through the latent instance space analysis, we now obtain a 2-dimensional latent instance space $\mathcal{H}$. To generate synthetic BBO problem instance from $\mathcal{H}$, several challenges still remain. One of them is how we model a proximal function space to represent or formulate mathematical functions of optimization problems as computer-understandable language. To address this, we adopt a symbolic tree representation for mathematical function, which has proven to be robust and effective in symbolic regression domain (Schmidt and Lipson 2009). To make it a comprehensive proximal space, we design a flexible symbol set containing diverse operators and operands, which facilitates accurate and efficient proximal representation. Another key challenge is how we reverse engineer a proximal representation of a BBO problem from the latent space $\mathcal{H}$. Genetic Programming (GP) (Koza 1992) is regarded as an effective way to discover functions from symbolic function spaces (Augusto and Barbosa 2000). In this paper, we propose a novel GP framework to search for the fittest function formula given a latent instance point in $\mathcal{H}$. This includes how we define the searching objective and how we enhance naive GP in both effectiveness and efficiency sides. We present the pseudocode of the GP search in Alg. 1 and elaborate our designs in the following sections.
@@
The pseudocode of the GP search in LSRE is provided in Alg. 1. In lines 2-3, we first randomly initialize a population of proximal tree representations $\{\tau_{i}\}_{i=1}^{N}$ by sampling from the symbolic space spanned by our proposed symbol set, and then evaluate them to attain their objective values $obj(\cdot)$. Then we begin an iterative searching process in lines 4-13. For each generation $g$, the solution population first undergoes a roulette_reproduction operation (line 6). It is a built-in reproduction strategy of gplearn, which randomly assign one of five default mutation/crossover operators for each tree individual, by a predetermined operator selection preference. After roulette_reproduction operation, an offspring population $\{\tau_{i}^{\prime}\}_{i=1}^{N}$ is obtained. In line 7, we propose a novel plug-and-play cross-dimensional local search strategy which can be easily integrated into the GP process to improve the searching accuracy. The motivation behind this local search strategy is that: although we could search for optimal function formulation, the dimension of the decision variables is not addressed. Several recent studies (Renau et al. 2020a; van Stein et al. 2023) reveal that even two function instances share mathematical formulation, their ELA features may vary due to different problem dimensions. To address this, in our proposed local search strategy, for each offspring $\tau_{i}^{\prime}$, we instantiate it to 2D, 5D and 10D function instances and compare their objective values, then the best objective value is regarded as the objective value of this offspring. We provide effectiveness validation in following ablation studies. After a total of $G$ generations GP searching, we obtain the best solution $\tau^{*}$ found so far (line 14) as well as its best-performing problem dimension.

### v2

@@ Abstract
To relieve intensive human-expertise required to design optimization algorithms, recent Meta-Black-Box Optimization (MetaBBO) researches leverage generalization strength of meta-learning to train neural network-based algorithm design policies over a predefined training problem set, which automates the adaptability of the low-level optimizers on unseen problem instances. Currently, a common training problem set choice in existing MetaBBOs is well-known benchmark suites CoCo-BBOB. Although such choice facilitates the MetaBBO’s development, problem instances in CoCo-BBOB are more or less limited in diversity, raising the risk of overfitting of MetaBBOs, which might further results in poor generalization. In this paper, we propose an instance generation approach, termed as LSRE, which could generate diverse training problem instances for MetaBBOs to learn more generalizable policies. LSRE first trains an autoencoder which maps high-dimensional problem features into a 2-dimensional latent space. Uniform-grid sampling in this latent space leads to hidden representations of problem instances with sufficient diversity. By leveraging a genetic-programming approach to search function formulas with minimal L2-distance to these hidden representations, LSRE reverse engineers a diversified problem set, termed as Diverse-BBO. We validate the effectiveness of LSRE by training various MetaBBOs on Diverse-BBO and observe their generalization performances on either synthetic or realistic scenarios. Extensive experimental results underscore the superiority of Diverse-BBO to existing training set choices in MetaBBOs. Further ablation studies not only demonstrate the effectiveness of design choices in LSRE, but also reveal interesting insights on instance diversity and MetaBBO’s generalization. We provide the code of LSRE and Diverse-BBO at https://github.com/MetaEvo/Diverse-BBO.
@@ 1 Introduction
Black-Box Optimization (BBO) problems, with neither the objective function nor corresponding derivative is accessible, are challenging to solve. Although Evolutionary Computation (EC) methods, such as Genetic Algorithm (GA), Particle Swarm Optimization (PSO), Differential Evolution (DE), as well as their modern (self-)adaptive variants are widely recognized as effective solutions, they often require tuning or even re-design for novel scenario adaption.
@@
To automate such manual algorithm design tasks, recent Meta-Black-Box Optimization (MetaBBO) researches propose learning generalizable design policy through meta-learning (metabox; ma2025toward). MetaBBO naturally forms a bi-level algorithmic structure. At the meta level, a learnable algorithm design policy (e.g., a neural network) outputs flexible design choices (e.g., parameter control (GLEET; ConfigX) and algorithm selection (RLDAS)) for low-level optimizer. At the lower level, a BBO optimizer is deployed to optimize target problem with the given design choice. The meta-objective is to train a design policy that enhances low-level BBO process. Since such paradigm provides a promising way to learn automated optimization algorithm design, MetaBBO has been widely explored and discussed in recent two years (rl-surrogate; DesignX; kl-lsgo; huang2025evaluation; metade; mo2025autosgnn).
@@
Given that MetaBBO essentially follows meta-learning paradigm (meta-learning-1; meta-learning), it unavoidably requires domain generalization designs (meta-dataset-1), i.e., meta-augmentation of training data (meta-dataset), so as to ensure the adaption during the testing time. To be specific, as shown in the yellow box of Fig. 1, existing MetaBBO approaches typically train their policies over a training problem set. The diversity of this set assures that the meta-level policy generalizes across diverse optimization problem characteristics (neurela). Common choices of the problem set in existing MetaBBO approaches are synthetic benchmarks such as CoCo-BBOB (bbob2010). Although these representative problem instances hold a long-standing history and reputation in the evaluation of BBO methods, they are primarily hand-crafted to differentiate algorithms apart from each other in terms of optimization performance (benchmark-limitation). Recent studies (ma2025toward; metabox-v2) reveal that such design bias may not keeps pace with the training&evaluation objectives of MetaBBO, i.e., generalization toward unseen problems. One possible solution is deriving realistic instances directly from real-world application (realistic-bench). However, collecting and formulating realistic problems require deep expertise. More importantly, the corresponding expensive simulation costs makes this solution impractical (GP-BBOB-1). A more promising way is generating synthetic problem under proper diversity measurement (GP-BBOB-3; ma-bbob), however, to the best of our knowledge, related researches are still very limited.
@@
To address the aforementioned limitations, we propose a novel generation framework that discovers diverse synthetic optimization problems through latent space reverse engineering, termed as LSRE, to advance the benchmark development in BBO and MetaBBO fields. Specifically, LSRE comprises following key designs: 1) an instance space analysis (isa-smith2023instance) method that trains an autoencoder to map the Exploratory Landscape Analysis (ELA) (ela) features of problem instances into 2-dimensional latent space; 2) an expressive symbol set including operands and operators that spans a comprehensive functional instance space; 3) a genetic programming method that effectively searches for function instance within the latent space, with an enhanced local search strategy and parallel acceleration. By grid-sampling 256 instance points from the latent space, we repeatedly call LSRE to search optimal function formula for each of them, resulting in a collection of 256 synthetic problem instances Diverse-BBO with adequate diversity. We validate the diversity advantage of our method against existing baseline benchmarks through large-scale MetaBBO benchmarking. The comprehensive experimental results demonstrate the effectiveness of our LSRE framework and the superiority of the generated Diverse-BBO to baseline benchmarks. We now summarize the core contributions of this paper:
@@ 2 Related Works
Existing MetaBBO approaches can be divided into several lines according to their specific algorithm design tasks: 1) Algorithm Selection (AS), where predefined algorithms/operators are selected by the meta-level policy to solve given problems. Early researches explored selecting operators within DE frameworks (DEDDQN; DEDQN), and more recent approaches have investigated dynamic switching between multiple advanced optimizers along the optimization process (RLDAS); 2) Algorithm Configuration (AC), where given an existing optimizer, the meta-level policy dictates desired parameter values dynamically along the optimization process (LDE; metabbo-ne-2; GLEET; ConfigX); 3) Algorithm Generation (AG), where the meta-level policy constructs both individual components and the entire workflow to synthesize novel BBO algorithms through algorithm workflow composition (ALDes), novel update rules generation (Symbol) and language modelling (LLaMaCo; eoh; metabbo-icl-5). In addition, sutdies on MetaBBO have also been extended to automated feature learning (neurela; RLDE-AFL), efficient offline learning (q-mamba; Surr-RLDE) and direct solution manipulation (GLHF; B2OPT).
@@
Existing MetaBBO methods typically adopt synthetic problems from well-established BBO benchmarks (eg., CoCo-BBOB (bbob2010), IOHProfiler (iohprofiler), IEEE CEC Competition series (cec2021)) and platforms (eg., MetaBox-v1/v2 (metabox; metabox-v2)). While these benchmark functions have been widely recognized and used for evaluating BBO methods, they are primarily hand-crafted and are collected with inherent biases (benchmark-limitation-2), aimed at creating performance distinctions between optimization algorithms (benchmark-limitation). The existing benchmarks’ design bias and lack of diversity may constrain the MetaBBO to learn generalizable design policies that excel at not only the training problem distribution but also unseen problems with various landscape features (ma2025toward; neurela). Constructing and collecting realistic problems (realistic-bench; proteindocking) could be a potential solution, but it requires deep expertise and is further limited by simulation costs (GP-BBOB-1), ultimately rendering the solution impractical.
Generating diverse synthetic problems might be more promising, yet several challenges persist: 1) The random function generator (GP-BBOB-3) lacks control over the characteristics and diversity of generated problems; 2) A recent line of works explore the possibility of augmenting new problem instances through affine combination of existing instances in BBO benchmark (ig&dim-reduction-dietrich2022increasing). A representative method is MA-BBOB (ma-bbob), which claimed achieved efficient and fine-grained problem generation. However, a key limitation is such approaches might restrict themselves within the selected composition functions. 3) In the light of symbolic regression (SR), some exploratory studies leveraged symbolic discovery tool such as Genetic Programming (GP) (GP-SR) to generate diversified instances by searching corresponding mathematical symbolic formulas (GP-BBOB-1; GP-BBOB-2). However, such researches are still limited. More importantly, given the inherent expensive running of GP and the design bottlenecks in these methods, their actual generation effectiveness is still unsatisfactory. In this paper, we propose LSRE to address the limitations in the third line of work in both effectiveness and efficiency side. On the one hand, we incorporate LSRE with several novel designs: 1) Effective latent instance space analysis through autoencoder; 2) Expressive symbol set enabling both efficient generation and flexible operations; 3) Cross-dimensional local search strategy to further enhance the GP searching. On the other hand, we introduce high-performance distributed acceleration techniques to facilitate efficient GP running.
@@ 3.1 Latent Instance Space Analysis
To align with practical usage of potential users, one of the most important issues in our framework is to pre-define a target problem range $\mathcal{D}$ as the basis for constructing corresponding the latent instance space. While interests of different users may vary, in this paper we primarily select CoCo-BBOB (bbob2010) to serve as $\mathcal{D}$ due to its ease of acquisition and long-standing reputation within BBO evaluation domain. We have to note that our framework can be easily extended to other user-specific problem ranges. We leave step-by-step instructions for such cases at Appendix A.1111“Appendix.pdf” in our codebase.. Given the 24 basic functions ($f_{1}\sim f_{24}$) in CoCo-BBOB, we first augment these functions with five different problem dimensions $[2,5,10,30,50]$, resulting in 120 functions. Next, we further generate a sufficient number of instances by applying random rotation and shift transformations to these 120 functions:
@@
We construct latent instance space of $\mathcal{D}$ by combining the Exploratory Landscape Analysis (ELA) (ela) and popular dimension reduction strategy autoencoder (AE-0). We illustrate the workflow of such latent instance space analysis in Fig. 2, and briefly introduce how these two techniques cooperate with each other:
ELA. It is a widely used tool for instance space analysis (bbob-ela-analysis&dim-reduction-vskvorc2020understanding) of BBO problems. Groups of analytic or statistical features, such as convexity, global structure and meta-model fitting are computed in ELA, which has been proven informative in algorithm selection (ela4as). Given a BBO problem instance $f\in\mathcal{D}$, we obtain its ELA features vector $E_{f}$ using open-source ELA package: pflacco [^2] https://github.com/Reiyan/pflacco. Specifically, we select a subset of representative and computationally efficient features as $E_{f}$. The selection standard and concrete feature list are provided in Appendix A.2 .
@@
We leave the settings of the autoencoder such as its architecture and corresponding training protocol in Appendix A.3. Once the training of the autoencoder finishes, we could obtain latent representation $h\in\mathbb{R}^{2}$ of a BBO problem instance through the encoder. More importantly, the well-trained model allows us to measure the diversity of problem instances in a 2-dimensional space, which further facilitates uniform instance sampling from this hidden space, since directly sampling from high-dimensional ELA space is neither efficient nor practical. It is worth mentioning that principal component analysis (PCA) (pca) is also a popular dimension reduction approach widely adopted in machine learning community. The reason behind our autoencoder choice is that: 1) PCA shows effectiveness in cases where features hold linear correlation assumption (AE-1; AE), however, different feature groups in ELA are independently computed; 2) We have conducted a preliminary study on using PCA for ELA feature reduction, the results show that the first two principle components of certain portion of instances in $\mathcal{D}$ can not surpass 51% importance threshold. We also provide an ablation study in Sec. 4.3.
@@ 3.2 Genetic Programming Search
Through the latent instance space analysis, we now obtain a 2-dimensional latent instance space $\mathcal{H}$. To generate synthetic BBO problem instance from $\mathcal{H}$, several challenges still remain. One of them is how we model a proximal function space to represent or formulate mathematical functions of optimization problems as computer-understandable language. To address this, we adopt a symbolic tree representation for mathematical function, which has proven to be robust and effective in symbolic regression domain (SR). To make it a comprehensive proximal space, we design a flexible symbol set containing diverse operators and operands, which facilitates accurate and efficient proximal representation. Another key challenge is how we reverse engineer a proximal representation of a BBO problem from the latent space $\mathcal{H}$. Genetic Programming (GP) (GP-SR) is regarded as an effective way to discover functions from symbolic function spaces (GP-SR-1). In this paper, we propose a novel GP framework to search for the fittest function formula given a latent instance point in $\mathcal{H}$. This includes how we define the searching objective and how we enhance naive GP in both effectiveness and efficiency sides. We present the pseudocode of the GP search in Alg. 1 and elaborate our designs in the following sections.
@@
The pseudocode of the GP search in LSRE is provided in Alg. 1. In lines 2-3, we first randomly initialize a population of proximal tree representations $\{\tau_{i}\}_{i=1}^{N}$ by sampling from the symbolic space spanned by our proposed symbol set, and then evaluate them to attain their objective values $obj(\cdot)$. Then we begin an iterative searching process in lines 4-13. For each generation $g$, the solution population first undergoes a roulette_reproduction operation (line 6). It is a built-in reproduction strategy of gplearn, which randomly assign one of five default mutation/crossover operators for each tree individual, by a predetermined operator selection preference. After roulette_reproduction operation, an offspring population $\{\tau_{i}^{\prime}\}_{i=1}^{N}$ is obtained. In line 7, we propose a novel plug-and-play cross-dimensional local search strategy which can be easily integrated into the GP process to improve the searching accuracy. The motivation behind this local search strategy is that: although we could search for optimal function formulation, the dimension of the decision variables is not addressed. Several recent studies (dimension-ela-1; dimension-ela-2) reveal that even two function instances share mathematical formulation, their ELA features may vary due to different problem dimensions. To address this, in our proposed local search strategy, for each offspring $\tau_{i}\prime$, we instantiate it to 2D, 5D and 10D function instances and compare their objective values, then the best objective value is regarded as the objective value of this offspring. We provide effectiveness validation in following ablation studies. After a total of $G$ generations GP searching, we obtain the best solution $\tau^{*}$ found so far (line 14) as well as its best-performing problem dimension.